Waterbury Insurance Company wants to study the relationship between the amount of fire damage and the distance between the bua-2. Is there a direct or indirect relationship between the distance from the fire station and the amount of fire damage? TheC-1. Determine and interpret the coefficient of determination. (Round your answer to 3 decimal places.) Coefficient of determd-1. Determine the correlation coefficient. (Round your answer to 3 decimal places.) Correlation coefficient d-2. Choose thed-3. How did you determine the sign of the correlation coefficient? It is because the slope is e-1. State the decision rule fe-2. Compute the value of the test statistic. (Round your answer to 2 decimal places.) Value of the test statistic e-3. Is th
Answer #1

a-1

Here the regression line

y = 1381345 + 300415 x

a-2

Here the distance between fire station and the amount of fire damage have a direct relationship.

b. Here x = 7 miles

y = 1381345 + 300415 * 7 = 3484250

c-1

Here coefficient of determination = r2 = SSR/SST = 1805.5782/3081.0716 = 0.586

c-2

Here 58.60% in the variation in damage is explained by the variation in distance.

d-1

Here the correlation coefficient = r = sqrt(0.586) = 0.7655

d-2

There is fairly solid positive link between the variables.

d-3 Here the coefficient of slope is positive so coefficient is positive.

e-1 tcritical = TINV(0.01, 28) = 2.7633

e-2 test statistic

T = 6.3

e-3 Here we will reject the null hypothesis if t < -2.7633 and t > 2.7633